OpenStudy (decarr432):

the domain of the function f(x)= (3x-6)/(x)+4.5 is {-3,-1,2,4,5} the range for the given f(x)= (3x-6)/(x)+4.5 is ?

OpenStudy (decarr432):

@Loser66

OpenStudy (decarr432):

Would this be set up as last time loser?

OpenStudy (loser66):

yup

OpenStudy (decarr432):

okay so

OpenStudy (decarr432):

3(-3)-6=-12/-3=4

OpenStudy (loser66):

3*(-3) =-9 -9-6 =-12???

OpenStudy (decarr432):

whoa my bad lol -15 haha

OpenStudy (loser66):

yes, then /-3= 5 then+4.5 =?

OpenStudy (decarr432):

so 4.5

OpenStudy (decarr432):

would equal

OpenStudy (decarr432):

-10.5

OpenStudy (loser66):

Please, type your expression carefully is it : \(f(x) = \dfrac{3x-6}{x} +4.5\)

OpenStudy (decarr432):

Yeah

OpenStudy (loser66):

ok, so, for the first one, x =-3, you get 5+4.5 = 10.5???

OpenStudy (decarr432):

so would i divide then add

OpenStudy (loser66):

\(f(-3) =\dfrac{3*(-3) -6}{-3}+4.5 = \dfrac{-15}{-3}+4.5 = 5+4.5 =?\)

OpenStudy (decarr432):

9.5 right

OpenStudy (loser66):

yup do the same with others.

OpenStudy (decarr432):

okay so

OpenStudy (decarr432):

\[3(-1)-6= -9\]

OpenStudy (decarr432):

\[\frac{ -9 }{ -1 }+4.5=9+4.5=13.5\]